OpenStudy (anonymous):

Solve 8^(x - 2) = 3 A) x ? 2.52 B) x ? 0.26 C) x ? 2.74 D) x ? 0.40

OpenStudy (anonymous):

take the natural log of both sides and isolate x

OpenStudy (anonymous):

1. Solve 8^(x-2) = 3. Take the logarithm of each side: log[8^(x-2)] = log[3]. Use the properties of logarithms to rewrite this as (x-2)log[8] = log[3]. Divide both sides by log[8], giving x - 2 = log[3]/log[8]. Add 2 to both sides, giving x = log[3]/log[8] + 2, which is approximately 2.52 (choice a).

OpenStudy (anonymous):

Hope that sums it up